# A Heuristic Philosophical Discourse on Various Applications of

### v1

@@
Abstract Differential Geometry in Quantum Gravity Researchthanks: This paper is a philosophical distillation of the basic concepts and central results of this author’s ongoing research project, spanning the last three decades, of applying Anastasios Mallios’s Abstract Differential Geometry to the ‘persistently stubborn’ problem of formulating a conceptually sound, mathematically consistent and calculationally finite Quantum Theory of Gravity. The paper will be posted at the General Relativity and Quantum Cosmology website www.arXiv.org/gr-qc before October 2024. This work may be viewed as the philosophical résumé and aftermath (:after the Maths!) of the following published papers [62, 63, 64, 73, 74, 75, 76, 77, 78, 79, 84, 85], as well as a sequel to the Dodecalogue paper [79] in the aftermath of the recent publication [80] and the pre-prints [81, 82, 83], which are currently work-in-progress in the pipeline. In turn, a longer version of the paper will constitute a chapter in a research monograph type of book that we had and have been working on, in collaboration with the late Professor Anastasios Mallios, since 2003 [65].
@@
(Thursday, 15th of August 2024
(Version I))
@@ Abstract
In the present paper, we outline and expound the fundamental and novel qualitative-cum-philosophical premises, principles, ideas, concepts, constructions and results that originate from our ongoing research project of applying the conceptual panoply and the technical machinery of Abstract Differential Geometry (ADG) to various persistently outstanding issues in Quantum Gravity (QG) [62, 63, 64, 73, 74, 75, 76, 77, 78, 79, 84, 85]. This paper may be regarded as a sequel to the paper [79] in the aftermath of the paper [80], which is currently in press. At the end of the paper, we discuss the potential philosophical repercussions of two possible future research routes that the main stream of our applications of ADG to QG may bifurcate towards in view of three independent, but overlapping, research papers that are currently under development [81, 82, 83].
@@
Key words: Natural Philosophy of Quantum Gravity and Quantum Yang-Mills Gauge Theories; Abstract Differential Geometry; sheaf theory; sheaf cohomology; category theory; topos theory; geometric prequantisation; canonical quantisation; background spacetime manifold independence; quantum gravity as a purely quantum gauge theory; gauge theory of the third kind; third sheaf cohomological quantisation of gravity and gauge theories
@@ 1 Prolegomena-cum-Motivation: Why Adopt a Philosophical Stance in Quantum Gravity Research?
There is currently a plethora of various and glaringly diverse approaches to QG, and it is not the ‘vain’ or ‘quixotic’ aim of this paper to list them all herein, let alone to review their conceptual and technical import or their successes and shortcomings. On the one hand, there is no unanimous agreement on what the ‘right’ or ‘correct’ approach to QG is and the diversity of the various different approaches—from string theory [25, 91], [^2] For current developments in this field, see the second reference above [91]. to loop quantum gravity [2, 3, 4, 44] [^3] For an exhaustive modern exposition of loop quantum gravity, see the last reference [44] above. and causal sets [6, 93, 94, 95], for example—exemplifies exactly that.
@@
Expository Declaration 1: In the sections that follow, whenever we discuss and analyse a qualitative-philosophical aspect of ADG-gravity, at the end of the discussion we will mark it by boldface markers in brackets like, for example, (F1): this would mean that the qualitative/philosophical characteristic of ADG-gravity being analysed and discussed satisfies the Finiteness 1 aspect of the EMF1 triplet above.
@@
Expository Declaration 2: It is a conscious decision and choice of this author not to include a single quantitative expression (e.g., technical mathematical formula) in the present paper, which is purely of a qualitative (discursive) character. For concise definitions of formal technical concepts and their associated mathematical formulae/modelling/equations, the reader will be directly referred to the relevant published literature.
Expository Declaration 3: Our philosophical treatise below is organised in short sections that end with ‘Aphorisms’—short ‘statements’ that distill the main philosophical gist of each section.
@@
$\bullet$ In turn, the differential manifold $M$, as a geometrical point-set, is equivalent to the structure algebra sheaf $\mathcal{C}^{\infty}(M)$ of germs of smooth coordinate functions of its points.
$\bullet$ In GR, the dynamical law of gravity is formulated as Einstein’s nonlinear partial differential equations of a $\mathcal{C}^{\infty}$-smooth spacetime metric $g_{\mu\nu}$ (and its derivatives), whose $10$ components are supposed to represent the gravitational field potentials.
@@
Fundamental Vicious Circle. If we wish to formulate the dynamical laws of QG as differential equations proper, it seems that we have to use the concepts and techniques of CDG on a smooth manifold. However, the latter is responsible for both the singularities of GR and the unphysical the infinities of QFT—sites in the spacetime manifold where the laws of physics appear to break down or lead to unphysical infinities for important observable quantities; hence, we seem to arrive at an impasse. [^8] The reader should note here that, in the three quotes above, both Feynman and Isham question the Mathematical Efficacy of CDG in QG (M1,F1, M2).
@@ 3 Enter ADG
Below, we itemise the basic tenets of ADG:
$\bullet$ Abstract Differential Geometry is more of a Leibnizian (:relational), rather than Newtonian (:geometrical), purely algebraic (:sheaf-theoretic and homological algebraic) way of doing differential geometry (:Calculus) [15, 8], without at all recourse to or dependence on a pointed, smooth geometric locally Euclidean background space (:a $\mathcal{C}^{\infty}$-smooth manifold) for its concepts, technical machinery and constructions thereof [47, 48, 53].(M2)
$\bullet$ The most fundamental concept of ADG is that of connection (viz. generalised differential) $\mathcal{D}$ acting on a vector sheaf $\mathbf{\mathcal{E}}$ over a suitably algebraized, by a certain so-called algebra structure sheaf $\mathbf{A}$ of generalised coordinates or what Mallios coined ‘arithmetics, arbitrary topological space $X$ as a linear and Leibnizian sheaf morphism. The pair $(\mathcal{D},\mathbf{\mathcal{E}})$ is coined an ADG-field.(M1,M2)
$\bullet$ Based on the concept of connection, ADG erects the whole edifice of CDG (plus more), but in the manifest absence of a background geometrical manifold.
$\bullet$ Thus, a new, entirely algebraic (:relational) notion of geometry emerges, whereby, geometry does not pertain to the configuration states (‘shape’) and measurement of objects living in an a priori posited (:postulated) ether-like background space [20], but rather, it derives from the algebraic (:dynamical) relations between the objects that live on that ‘space’. [^9] These ‘objects’ are the very ADG-connection fields acting on the sections of the vector sheaves involved.
$\bullet$ The last bullet point put in more physical terms: physical geometry (:physical ‘spacetime’) is not a priori posited like the differential spacetime manifold $M$ of GR. Rather, it derives from the dynamics (:the differential equations involving the connection field $\mathcal{D}$) of the ‘objects’ (:the dynamical physical fields themselves).(E2,M2)
@@
Aphorism 3. From the ADG-theoretic perspective, physical geometry (or physical ‘spacetime’) derives from, or is the outcome of, the algebraic dynamical relations between the ADG-fields (:the physical laws, which are formulated categorically as equations between he relevant sheaf morphisms that the ADG-connection fields correspond to). [^10] In this regard, one may think of the more commonly used mathematical term ‘solution space’ derived from a set of (differential) equations set up (on a manifold) by the usual CDG-means. That ‘solution space’ is the ‘physical geometry’.(E2,M2)
@@
$\bullet$ In both the point of pointlessness and finiteness, ADG has been applied towards formulating on the one hand a locally finite, causal and quantal version of Lorentzian vacuum Einstein gravity and free Yang-Mills theories, and on the other, the same dynamical equations are seen to hold over spaces that are everywhere dense with singularities of the most unmanageable kind from the point of view of ADG [50, 51, 55, 56, 62, 63, 64, 66, 67, 68, 75].
@@
Aphorism 4. We can formulate the dynamical equations of Einstein and Yang-Mills over highly pathological and problematic spaces, especially when viewed from the smooth background spacetime manifold perspective of CDG. Thus, singularities (and their associated infinities) are not insuperable obstacles and ‘sites’ where the differential equations that represent the dynamical field laws of Nature appear to break down. Not only we can evade them by ADG-theoretic means, but also we can ‘calculate’ (:do Calculus!) in their very presence, in spite of them. The inherently algebraic differential geometric mechanism of ADG is genuinely background smooth spacetime independent, hence it does stumble on its inherent anomalies and pathologies [50, 51, 55, 56, 62, 63, 64, 66, 67, 68, 75]. (E1,M1,F1,E2,M2)
@@ 4.1 ADG Gravity: Einstein’s Purely Algebraic Description of Reality in the Quantum Deep
In the Philosophy of Physics there is a well established view that Einstein in vain pursued his Unified Field Theory [^12] Einstein had originally coined it Unitary Field Theory instead [17]. research on a continuous spacetime manifold in spite of the inherently finitistic and algebraic description of Physical Reality at subatomic scales that Quantum Theory brought about.
@@
In connection with the two Einstein quotes (Q5,6) above, with our fifth Aphorism next, which closes this section, we kill two birds with one stone:
Aphorism 5. ADG is an entirely algebraic method for formulating gravity and quantum Yang-Mills theories of matter field theoretically and finitistically by evading singularities and without any dependence on a background geometrical spacetime continuum, with its all inherent singularities and associated unphysical smooth field infinities [64, 75, 78, 79, 80].
@@ 4.2 Sheaf Theory and the Transition from Local to Global
Structures closely related to sheaves are special type of categories called topoi [46, 45], which are pointless spaces having their own internal logic. Topoi have been applied to both quantum logic [9, 10, 11, 12] and quantum spacetime structures, including QG research [35, 36, 37, 38, 39, 40, 76, 77, 80, 69, 70]
@@ 5 Revisiting Feynman: Gravity as Gauge Theory
Returning to the Feynman quote in Section 2, we wish to dwell a bit on his remark that the fact that the gravitational field was identified with the smooth spacetime metric $g_{\mu\nu}$ of the CDG-based Riemannian geometry on a differential spacetime manifold, was an ‘accident’ of theory making. [^19] That is to say, Einstein formulated GR as the dynamics of the metric, ‘because’ he used the CDG-based Riemannian geometry of a smooth base spacetime manifold [71]. Rather, Feynman intuited that:
The deeper character of gravity is that it is a gauge force, much like the other three fundamental forces, while the methods of CDG would be ineffective in the QG deep. [^20] Similarly to what Isham said in the quote following Feynman’s.
@@ 5.1 Enter ADG: Gravity as Pure Gauge Theory of the Third Kind
Aphorism 6: Gravity is a pure gauge theory, without recourse to an underlying (smooth) spacetime manifold structure for either its mathematical formulation or its physical interpretation. Gravity involves the dynamics of the ADG-theoretic Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, which is simply a connection $\mathcal{D}$ on a suitable vector sheaf $\mathbf{\mathcal{E}}$. The dynamical Einstein equations are derived from a variational principle applied on the ADG-version of the Einstein-Hilbert action functional [53, 64, 80]. [^23] By ‘pure gauge theory’ above, it is meant that the sole dynamical variable in the theory is the gravitational connection $\mathcal{D}$, acting on the sections of a suitable vector sheaf $\mathbf{\mathcal{E}}$, and nothing else. The corresponding formalism has been coined half-order formalism, to distinguish it from the first-order formalism of Ashtekar and its Loop QG outgrowth, and of course from the usual second-order formalism of the original Einstein theory (GR). (E1,M1,F1)
@@
Space and time are human inventions convenient for representing, localising and quantifying our measurements of physical observable entities.
@@
By contrast, in ADG-gravity, where we have no background spacetime manifold and the ADG-field autodynamics is purely gauge,
@@
whereby, as noted earlier, $\mathbf{A}$ is the structure sheaf of generalised arithmetics or coordinates in the theory.
In other words, the ADG-gravitational dynamics, which is formulated entirely categorically in terms of the connection $\mathcal{D}$ sheaf morphism, is respected by (:‘remains invariant under’) all our generalised measurements (:arithmetics, event coordinate determinations) encoded in the structure sheaf $\mathbf{A}$. The gravitational dynamics ‘sees through’ all our coordinate measurements (:spacetime event localisations) in $\mathbf{A}$ [79, 80].

### v2

@@
Abstract Differential Geometry in Quantum Gravity Researchthanks: This paper is a philosophical distillation of the basic concepts and central results of this author’s ongoing research project, spanning the last three decades, of applying Anastasios Mallios’s Abstract Differential Geometry to the ‘persistently stubborn’ problem of formulating a conceptually sound, mathematically consistent and calculationally finite Quantum Theory of Gravity. The paper is posted at the General Relativity and Quantum Cosmology website www.arXiv.org/gr-qc, as follows: gr-qc/2408.07725. This work may be viewed as the philosophical résumé and aftermath (:after the Maths!) of the following published papers [67, 68, 69, 78, 79, 80, 81, 82, 83, 84, 89, 90], as well as a sequel to the Dodecalogue paper [84] in the aftermath of the recent publication [85] and the pre-prints [86, 87, 88], which are currently work-in-progress in the pipeline. In turn, a longer version of the paper will constitute a chapter in a research monograph type of book that we had and have been working on, in collaboration with the late Professor Anastasios Mallios, since 2003 [70].
@@
(Tuesday, 20th of August 2024
(Version II))
@@ Abstract
In the present paper, we outline and expound the fundamental and novel qualitative-cum-philosophical premises, principles, ideas, concepts, constructions and results that originate from our ongoing research project of applying the new conceptual panoply and the novel technical machinery of Abstract Differential Geometry (ADG) to various persistently outstanding issues in Quantum Gravity (QG) [67, 68, 69, 78, 79, 80, 81, 82, 83, 84, 89, 90]. This paper may be regarded as a sequel to the paper [84] in the aftermath of the paper [85], which is currently in press. At the end of the paper, we discuss the potential philosophical repercussions of two possible future research routes that the main stream of our applications of ADG to QG may bifurcate towards in view of three independent, but overlapping, research papers that are currently under development [86, 87, 88].
@@
Key words: Natural Philosophy of Quantum Gravity and Quantum Yang-Mills Gauge Theories; Abstract Differential Geometry; Sheaf Theory; Sheaf Cohomology; Category Theory; Topos Theory; Geometric Prequantisation; Canonical Quantisation; background spacetime manifold independence; Quantum Gravity as a purely Quantum Gauge Theory; gauge theory of the third kind; sheaf cohomological third quantisation of gravity and gauge theories
@@ 1 Prolegomena-cum-Motivation: Why Adopt a Philosophical Stance in Quantum Gravity Research?
There is currently a plethora of various and glaringly diverse approaches to QG, and it is not the ‘vain’ or ‘quixotic’ aim of this paper to list them all herein, let alone to review their conceptual and technical import or their successes and shortcomings. On the one hand, there is no unanimous agreement on what the ‘right’ or ‘correct’ approach to QG is and the diversity of the various different approaches—from string theory [24, 96], [^2] For current developments in this field, see the second reference above [96]. to loop quantum gravity [2, 3, 4, 49] [^3] For an exhaustive modern exposition of loop quantum gravity, see the last reference [49] above. and the ‘discrete’ causal sets [6, 97, 98, 99], for example—exemplifies exactly that.
@@
Expository Declaration 1: In the sections that follow, whenever we discuss and analyse a qualitative-philosophical aspect of ADG-gravity, at the end of the discussion we will mark it by boldface markers in brackets like, for example, (F1): this would mean that the qualitative/philosophical characteristic of ADG-gravity being analysed and discussed satisfies the Finiteness 1 aspect of the EMF1 triplet outlined above.
@@
Expository Declaration 2: It is a conscious decision and choice of this author not to include a single quantitative expression (e.g., technical mathematical formula) in the present paper, which is purely of a qualitative (discursive) character. For concise definitions of formal technical concepts and their associated mathematical formulae/modelling/equations and associated calculations using ADG, the reader will be directly referred to the relevant published literature.
Expository Declaration 3: Our philosophical treatise below is organised in short sections that end with ‘Aphorisms’—short and terse statements that distill the main philosophical gist of each section.
@@
$\bullet$ In turn, the differential manifold $M$, as a geometrical point-set, is equivalent to the structure algebra sheaf $\mathcal{C}^{\infty}(M)$ of germs of smooth coordinate functions of its points. [^6] This is an example of the general notion of Gelfand Duality that is one of the fundamental pillars on which ADG was originally erected by Mallios [52, 53, 58] and used extensively in its applications to QG thereafter [67, 68, 69, 80, 81, 83, 84], and very recently in [85].
$\bullet$ In GR, the dynamical law of gravity is formulated as Einstein’s nonlinear partial differential equations of a $\mathcal{C}^{\infty}$-smooth spacetime metric $g_{\mu\nu}$ (and its derivatives), whose $10$ components are supposed to represent the smooth gravitational field potentials.
@@
Fundamental Vicious Circle. If we wish to formulate the dynamical laws of QG as differential equations proper, it seems that we have to use the concepts and techniques of CDG on a smooth manifold. However, the latter is responsible for both the singularities of GR and the unphysical the infinities of QFT—sites in the spacetime manifold where the laws of physics appear to break down or lead to unphysical infinities for important observable field quantities; hence, we seem to arrive at an impasse. [^9] The reader should note here that, in the three quotes above, both Feynman and Isham question the Mathematical Efficacy of CDG in QG (M1,F1,M2).
@@ 3 Enter ADG
Below, in juxtaposition and contradistinction to the points outlined above, we itemise the basic tenets of ADG:
$\bullet$ Abstract Differential Geometry is more of a Leibnizian (:relational), rather than Newtonian (:geometrical), purely algebraic (:sheaf-theoretic and homological algebraic) way of doing differential geometry (:Calculus) [14, 8], without at all recourse to or dependence on a pointed, smooth geometric locally Euclidean background spacetime (:a $\mathcal{C}^{\infty}$-smooth manifold) for its concepts, technical machinery and constructions thereof [52, 53, 58].(M2)
$\bullet$ The most fundamental concept of ADG is that of connection (viz. generalised differential) $\mathcal{D}$ acting on a vector sheaf $\mathbf{\mathcal{E}}$ over a suitably algebraized, by a certain so-called algebra structure sheaf $\mathbf{A}$ of generalised coordinates or what Mallios originally coined ‘arithmetics [52, 58], arbitrary topological space $X$ as a linear and Leibnizian sheaf morphism. The pair $(\mathcal{D},\mathbf{\mathcal{E}})$ is coined an ADG-field [52, 58].(M1,M2)
$\bullet$ Based on the concept of connection, ADG erects the whole edifice of CDG (plus more), but in the manifest absence of a background geometrical manifold [52, 58].
$\bullet$ Thus, a new, entirely algebraic (:relational) notion of geometry emerges, whereby, geometry does not pertain to the configuration states (‘shape’) and measurement of objects living in an a priori posited (:postulated) ether-like background space [19, 20], but rather, it derives from the algebraic (:dynamical) relations between the objects that live on that ‘space’. [^10] These ‘objects’ are the very ADG-connection fields acting on the sections of the vector sheaves involved.
$\bullet$ The last bullet point put in more physical terms: physical geometry (:physical ‘spacetime’) is not a priori posited like the differential spacetime manifold $M$ of GR. Rather, it derives from the dynamics (:the differential equations involving the connection field $\mathcal{D}$) of the ‘objects’ (:the dynamical physical fields) themselves. [^11] The reader should refer to the wonderful book [75] for a thorough technical exposition and semantic explanation of our ‘motto’ here, namely that: geometry and spacetime derive from the ADG-field dynamics, not the other way around. (E2,M2)
@@
Aphorism 3. From the ADG-theoretic perspective, physical geometry (or physical ‘spacetime’) derives from, or is the outcome of, the algebraically (:sheaf theoretically) represented dynamical relations between the ADG-fields (:the physical laws, which are formulated categorically as equations between he relevant sheaf morphisms that the ADG-connection fields correspond to). [^12] In this regard, one may think of the more commonly used mathematical term ‘solution space’ derived from a set of (differential) equations set up (on a manifold) by the usual CDG-means. That ‘solution space’ is the ‘physical geometry’ [75, 85].(E2,M2)
@@
$\bullet$ In both the point of pointlessness and finiteness, ADG has been applied towards formulating on the one hand a locally finite, causal and quantal version of Lorentzian vacuum Einstein gravity and free Yang-Mills theories, and on the other, the same dynamical equations are seen to hold over spaces that are everywhere dense with singularities of the most unmanageable and problematic kind from the point of view of CDG [55, 56, 60, 61, 67, 68, 69, 71, 72, 73, 80].
@@
Aphorism 4. We can formulate the dynamical equations of Einstein and Yang-Mills over highly pathological and problematic spaces, especially when viewed from the smooth background spacetime manifold perspective of CDG. Thus, singularities (and their associated infinities) are not insuperable obstacles and ‘sites’ where the differential equations that represent the dynamical field laws of Nature appear to break down. Not only we can evade them by ADG-theoretic means, but also we can ‘calculate’ (:actually do Calculus!) in their very presence, in spite of them. The inherently algebraic differential geometric mechanism of ADG is genuinely background smooth spacetime independent, hence it does stumble on its inherent anomalies and pathologies [55, 56, 60, 61, 67, 68, 69, 71, 72, 73, 80]. (E1,M1,F1,E2,M2)
@@ 4.1 ADG Gravity: Einstein’s Purely Algebraic Description of Reality in the Quantum Deep
In the Philosophy of Physics there is a well established view that Einstein stubbornly, but in vain, pursued his Unified Field Theory [^14] Einstein had originally coined it Unitary Field Theory instead [16]. research on a continuous spacetime manifold in spite of the inherently finitistic and algebraic description of Physical Reality at subatomic scales that Quantum Theory brought about.
@@
In connection with the two Einstein quotes (Q5,6) above, with our fifth Aphorism next, which closes this section, we kill two birds with one stone as it were:
Aphorism 5. ADG is an entirely algebraic method for formulating gravity and quantum Yang-Mills theories of matter field theoretically and finitistically by evading singularities, plus without any dependence on a background geometrical spacetime continuum, with its all inherent singularities and associated unphysical smooth field infinities [69, 80, 83, 84, 85].
@@ 4.2 Sheaf Theory and the Transition from Local to Global
Structures closely related to sheaves are special type of categories called topoi [51, 50], which are pointless spaces having their own internal, intuitionistic-type of logic. Topoi have been applied to both quantum logic [9, 10, 11, 12] and quantum spacetime structures, including QG research [34, 35, 36, 37, 38, 39, 81, 82, 85, 74, 75]. [^19] We also recommend that the reader refers to Goro Kato’s truly original application of sheaf and topos theory to quantum spacetime structure [41, 42, 43, 44, 45, 46].
@@ 5 Revisiting Feynman: Gravity as Gauge Theory
Returning to the Feynman quote in Section 2, we wish to dwell a bit on his remark that the fact that the gravitational field was identified with the smooth spacetime metric $g_{\mu\nu}$ of the CDG-based Riemannian geometry on a differential spacetime manifold, was an ‘accident’ of theory making. [^22] That is to say, Einstein formulated GR as the dynamics of the metric, ‘because’ he used the CDG-based pseudo-Riemannian geometry of a smooth base spacetime manifold [76]. Rather, Feynman intuited that:
The deeper character of gravity is that it is a gauge force, much like the other three fundamental forces, while the methods of CDG would be ineffective in the QG deep. [^23] Similarly to what Isham said in quote (Q4) following Feynman’s (Q2,3).
@@ 5.1 Enter ADG: Gravity as Pure Gauge Theory of the Third Kind
Aphorism 6: Gravity is a pure gauge theory, without recourse to an underlying (smooth) spacetime manifold structure for either its mathematical formulation or its physical interpretation. Gravity involves the dynamics of the ADG-theoretic Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, which is simply a connection $\mathcal{D}$ on a suitable vector sheaf $\mathbf{\mathcal{E}}$. The dynamical Einstein equations are derived from a variational principle applied on the ADG-version of the Einstein-Hilbert action functional [58, 69, 83, 85]. [^26] By ‘pure gauge theory’ above, it is meant that the sole dynamical variable in the theory is the gravitational connection $\mathcal{D}$, acting on the sections of a suitable vector sheaf $\mathbf{\mathcal{E}}$, and nothing else. The corresponding formalism has been coined half-order formalism, to distinguish it from the first-order formalism of Ashtekar and its smooth spin-Lorentzian connection holonomy based Loop QG outgrowth [49], and of course from the usual second-order formalism of the original Einstein theory (GR) [76]. (E1,M1,F1)
@@
Space and time are human inventions convenient for representing, localising and quantifying our measurements of physically observable entities.
@@
By contrast, in ADG-gravity, where we have no smooth background spacetime manifold and the ADG-field autodynamics is purely gauge,
@@
whereby, as noted earlier, $\mathbf{A}$ is the abelian (:commutative) algebra structure sheaf of generalised arithmetics or coordinates in the theory.
In other words, the ADG-gravitational dynamics, which is formulated entirely categorically in terms of the connection $\mathcal{D}$ sheaf morphism, is respected by (:‘remains invariant under’) all our generalised measurements (:arithmetics, event coordinate determinations) encoded in the abelian structure sheaf $\mathbf{A}$. The gravitational dynamics ‘sees through’ all our coordinate measurements (:spacetime event localisations) in $\mathbf{A}$ [84, 85].